### v3

# Magnitude homology is a derived functor

Abstract.
We prove that the magnitude (co)homology of an enriched category can, under some technical assumptions, be described in terms of derived functors between certain abelian categories. We show how this statement is specified for the cases of quasimetric spaces, finite quasimetric spaces, and finite digraphs. For quasimetric spaces, we define the notion of a distance module over a quasimetric space, define the functor of (co)invariants of a distance module and show that the magnitude (co)homology can be presented via its derived functors. As a corollary we obtain that the magnitude cohomology of a quasimetric space can be presented in terms of Ext functors in the category of distance modules. For finite quasimetric spaces, we show that magnitude (co)homology can be presented in terms of Tor and Ext functors over a certain graded algebra, called the distance algebra of the quasimetric space. For finite digraphs, the distance algebra is a bound quiver algebra. In addition, we show that the magnitude cohomology algebra of a finite quasimetric space can be described as a Yoneda algebra.
Contents
1. Introduction
In [leinster2008euler] Leinster introduced the notion of the Euler characteristic of a finite category. This Euler characteristic is a rational number coinciding with the Euler characteristic of the classifying space if it is a finite CW-complex. Later [leinster2013magnitude] he generalized this notion to the notion of the magnitude of a finite -category where is a monoidal category equipped with a multiplicative homomorphism to a ring . A special case of this notion is the notion of the magnitude of a finite metric space. The concept of magnitude was further explored in the special case of finite graphs [leinster2019magnitude], where the magnitude of a graph is defined as a formal power series with integer coefficients. It was shown that the coefficients of the formal power series can be described as a certain alternating sum, similar to the sum arising in the classical definition of the Euler characteristic. Having this analogy in mind, Hepworth and Willerton [hepworth2017categorifying] defined the magnitude homology of a graph and showed that the coefficients of the magnitude power series are alternating sums of ranks of the magnitude homology groups. Leinster and Shulman [leinster2021magnitude] generalized the notion of the magnitude homology from the case of a graph to the case of a small -category where is a semicartesian symmetric monoidal category equipped with a strong monoidal functor to an abelian symmetric monoidal category As in the case of magnitude, a special case was that of metric space. They showed that the relation between the magnitude and the magnitude homology can be generalized from the case of graphs to the case of finite metric spaces. The dual theory of magnitude cohomology was developed by Hepworth [hepworth2022magnitude], where he introduces a product in the magnitude cohomology and shows that in the case of a finite metric space the magnitude cohomology algebra completely determines the metric space.
If is the category of sets, is the category of abelian groups and is the functor of free abelian group, then a -category is just a category, and the magnitude homology of a category coincides with its ordinary homology (homology of the classifying space). It is well known that the ordinary homology of a category can be represented as derived functors of the functor of a colimit [gabriel2012calculus, App.II,Prop.3.3], [quillen2006higher, §1]. To be more precise, if we denote by the category of functors to the category of abelian groups then we obtain a functor of colimit which is a right exact functor between abelian categories, and the homology of can be presented via it’s left derived functors
There is a similar statement about the cohomology of and the functor of limit. The goal of this work is to generalize these statements to the magnitude (co)homology of an enriched category, and point out the specifics that arise in the special cases of quasimetric space, finite quasimetric space and digraph. We also study the magnitude cohomology algebra of a finite quasimetric space and show that it can be presented as a Yoneda algebra.
In order to prove our main statement about the magnitude homology of an enriched category, we assume that is an abelian Bénabou cosmos and develop the theory of -modules where is a small -category. The category of -modules is defined as the category of -functors . We show that is an abelian category with enough of projectives and injectives, assuming that has enough of projectives and injectives. By an augmented -category we mean a small -category equipped with a functor into the single-object -category Similar to how this is done in the group homology theory, for an augmented -category we define functor of trivial module, and its left and right adjoint functors of coinvariants and invariants
and define the homology and cohomology of via their derived functors
We prove that, under some additional assumptions on this homology and cohomology can be computed using some version of the bar resolution.
Returning to the setting of the magnitude homology of enriched categories, we assume that is a Bénabou cosmos, is an abelian Bénabou cosmos and is a strong monoidal functor. Then for an augmented -category , satisfying some technical conditions, we prove that the magnitude (co)homology of can be presented as (co)homology of the corresponding -category obtained from using the cosmos change construction
Let us note that for our convenience here we have slightly generalized the definition of the magnitude homology, and abandoned the assumption that the monoidal category is semicartesian, replacing this assumption with the assumption that an augmentation is given on .
We also studied how these concepts are specialized for the case of quasimetric spaces. For a quasimetric space and a commutative ring we define the notion of a distance module over Roughly speaking, a distance module over consists of a collection of -graded -modules indexed by points of equipped with some -linear maps between them, satisfying some axioms. We prove that the category of distance modules over is equivalent to the category -modules
Using this, we give a more explicit description of the magnitude homology of a quasimetric space in terms of derived functors. As a corollary we obtain a description of the magnitude cohomology with coefficients in in terms of Ext-functors in the abelian category of distance modules
where is the trivial distance module and is this distance module with a shifted grading.
If is a finite quasimetric space, using these descriptions we show that the magnitude (co)homology can be described in terms of bigraded Tor and Ext functors over some -graded algebra called the distance algebra of
Here denotes a -module described as a quotient of by some explicit ideal. Moreover, in this case the multiplication on the magnitude cohomology can be described in terms of the Yoneda product.
We also have some versions of all the statements about distance modules which are relative to some subgroup Namely, if is a subgroup of the additive group of real numbers and is a quasimetric space such that all finite distances lie in then we have a notion of -graded distance module, -graded algebra and so on. This is especially useful when and all finite distances are integers.
Specifying to an even more special case of a finite digraph treated as a quasimetric space, we find that the distance algebra can be presented as the quotient algebra of the path algebra
by an admissible ideal defined by some explicit set of relations. In this case the -module can be presented as where is the ideal generated by all arrows. Moreover in this case the category of -graded distance modules can be described as the category of graded representations of the bound quiver
2. Background from enriched category theory
In this section we collect some definitions and results from enriched category theory that can be found in [kelly1982basic], [eilenberg1966closed] and [lawvere1973metric].
2.1. Bénabou cosmos
Further in this section, unless otherwise stated, we denote by a Bénabou cosmos i.e. a complete, cocomplete symmetric closed monoidal category. The underlying category is denoted by The tensor product in is denoted by and the unit object is denoted by The inner hom-object is denoted by while the hom-set is denoted by
Since the functors and are adjoint, we have the unit and counit of the adjunction
This counit allows to define the composition law for the inner hom of
as the adjoint to the composition of counits (see [kelly1982basic, §1.6])
If is a -category, the hom-object is denoted by and the hom-set in is denoted by The underlying category of will be also denoted by
2.2. Ends
For -categories and we can consider their tensor product and the opposite category [kelly1982basic, §1.4].
Let be a -functor. We say that a collection of morphisms indexed by objects of is -dinatural (or extraordinary -natural), if the diagram
is commutative (see [kelly1982basic, §1.7]). The universal -dinatural transformation is called the end of , and its domain is denoted by
If is small and then the end exists and can be computed as an equaliser [kelly1982basic, §2.1]
where the morphisms are defined by the equations and and the morphisms and are double adjoint morphisms to and from the diagram (2.4).
The coend is defined dually, and if is small and it exists and can be presented as a coequaliser
2.3. Functor categories
Assume that and are -categories. For -functors a -natural transformation is defined as a collection of morphisms indexed by objects of such that the diagram
is commutative.
If is small, we can define the category of functors whose objects are -functors and morphisms are -natural transformations. This category has a natural -enrichment defined by the formula
So we obtain a -category of -functors . Here we note that the underlying hom-set coincides with the collection of -natural transformations . Moreover, if and are small, there is an equivalence of -categories [kelly1982basic, §2.3]
Lemma 2.1.
Let be a -natural transformation such that is an isomorphism for each object Then is an isomorphism.
Proof.
We need to check that the collection is a -natural transformation. Since is a -natural transformation, we have If we compose it with we obtain Therefore is a -natural transformation. ∎
2.4. Weighted limits
Let be a small -category and let be -functors. Then the -weighted (or indexed) limit of can be defined as [kelly1982basic, (3.7)]
So, in this case it exists. Moreover, in this case we see that it is -natural in and
More generally, if and are -functors, then the weighted limit is defined as an object of together with an isomorphism
which is -natural in If we take then the image of under the isomorphism is denoted by
and called the unit of the weighted limit . The isomorphism (2.12) is uniquely defined by
A -category is called -complete, if all weighted limits of functors from a small -category exist. In this case the definition of a weighted limit can be uniquely extended to a -functor [kelly1982basic, (3.11) and §1.10]
such that the isomorphism (2.12) is -natural in and
A -functor is said to preserve the limit if there is an isomorphism such that the diagram
is commutative [kelly1982basic, §3.2].
A weighted colimit in can be defined as a weighted limit in More precisely, if we have functors and then the weighted colimit is an object of together with an isomorphism
which is -natural in If we take then the image of under the isomorphism is denoted by
and called the counit of the weighted colimit . The preservation of weighted colimits is defined dually to the preservation of weighted limits.
The category is called -cocomplete, if all weighted colimits from small categories exist. In this case the definition of a weighted colimit can be uniquely extended to a -functor [kelly1982basic, (3.11) and §1.10]
such that the isomorphism (2.16) is -natural in and
A -category is called -bicomplete, if it is -complete and -cocomplete. For example, itself is -bicomplete [kelly1982basic, (3.74)]. Moreover, if is a small -category, the functor category is -bicomplete [kelly1982basic, Prop. 3.75]. In this case for each object we can consider the evaluation -functor [kelly1982basic, (2.17)]
This -functor preserves all weighted limits and colimits [kelly1982basic, (3.16)]. In particular, for a functor we have isomorphisms
which are -natural in and .
For any object of we can consider -functors and Then the enriched Yoneda lemma says that for any functor there are isomorphisms
-natural in and [kelly1982basic, (3.12)].
For a functor its end and coend can be presented as weighted limits and colimits [kelly1982basic, (3.59),(3.66)]
Therefore, if is -bicomplete and is small, the ends and coends of functors exist.
2.5. Conical limits
For a set and an object of we denote by the coproduct of copies of indexed by Let be a small ordinary category (over the category of sets). We denote by the free -category generated by [kelly1982basic, §2.5]. Its objects are objects of and hom-objects are If is a -category, a functor uniquely defines a -functor In particular, the functor that sends all objects to and all morphisms to defines a functor The conical limit of is defined by
The unit (2.12) in this case is defined by the cone
where
If then the ordinary limit of the functor is always its conical colimit [kelly1982basic, §3.8]
So we can say that ordinary limits in also have a stronger property of being conical limits in the enriched sense. Then the defining property of the weighted limit of a functor (2.12) can be rewritten as the existence of the following -natural isomorphism.
If the conical limit exists in , then the ordinary limit also exists in and it coincides with the conical limit [kelly1982basic, (3.53)].
However, the property of being a conical limit in is in general stronger than being an ordinary limit in For example, the product of two objects in is an object with a natural isomorphism But the corresponding conical limit in is an object with a -natural isomorphism On the other hand, if then for the ordinary product we have the isomorphism
Conical colimits are defined dually, and their defining property is
If then the ordinary colimits in are conical colimits in
In any -category a conical colimit is also an ordinary colimit in
If is a functor that preserves the conical limit then there is an isomorphism such that the diagram (2.15) is commutative. It follows that the diagram
is also commutative. A dual digramm is commutative for a functor that preserves a conical colimit.
Since, conical limits and colimits are particular cases of weighted limits and colimits, they exist in any -bicomplete category In particular, if is a small category, then conical limits and colimits exist in and the evaluation -functor
preserves them (2.19). In particular, for any functor there are isomorphisms
which are -natural in and natural in
2.6. Kan extensions
Let and be small -categories and let be a complete -category. Consider the functor categories and Then any -functor defines a composition -functor.
Moreover, there are two functors
which are adjoint to
in the following enriched sense [kelly1982basic, Th. 4.38, Th. 4.50].
The adjunction in the enriched sense implies the adjunction in the ordinary sense.
The functor is called the left Kan extension of with respect to and is called the right Kan extension of with respect to They can be computed on objects as follows [kelly1982basic, (4.18),(4.9)].
Moreover, if then they can be presented in terms of ends and coends [kelly1982basic, (4.24),(4.25)].
2.7. Cosmos change
Let and be two Bénabou cosmoi. A lax monoidal functor is a functor together with a morphism
and a natural transformation
satisfying some axioms [eilenberg1966closed, p.473]. The natural transformation defines a natural transformation
as the composition of the morhism adjoint to (2.43) and a morphism induced by the counit This morphism defines a structure of a closed functor on (see [eilenberg1966closed, Ch.I,§3 and Ch.II,Prop.4.3]).
For a lax monoidal functor and a -category one can define a -category (see [lawvere1973metric, p.150-151] and [eilenberg1966closed, Ch.I, Prop.6.1]). Its objects are objects of hom-objects are defined by the composition in is defined as the composition
while the identity morphism in is defined as the composition
Any -functor defines a -functor such that
2.8. Nerve of an augmented category
For ordinary small category there is a notion of nerve, which is a simplicial set associated with a category. For enriched categories it is more complicated. One option is to assume that the underlying monoidal category is semicartesian. In this case the construction of nerve is well defined. This happens because any category enriched over a semicartesian monoidal category has a unique augmentation. If we want to consider the nerve of a category enriched over any monoidal category, we need to equip it by an augmentation.
In this subsection we assume that is cocomplete symmetric monoidal category (not necessarily Bénabou cosmos). Denote by the -category with one object and hom given by The composition and identity of are defined as the standard isomorphism and the identity morphism An augmented -category is a small non-empty -category together with a -functor called augmentation.
The nerve of an augmented -category is a simplicial object in whose components are defined as
In particular, for we have The face maps for are defined as follows. For they are induced by compositions
The morphisms and are induced by the augmentation and The degeneracy maps are defined induced by the identity map We leave the verification of simplicial relations to the reader.
3. Modules over an enriched category
3.1. Abelian Bénabou cosmoi
A Bénabou cosmos is called abelian if is an abelian category satisfying Grothendieck axioms and The Grothendieck axioms say that a small (co)product of short exact sequences is a short exact sequence. Further in this section we assume that is an abelian Bénabou cosmos.
Proposition 3.1.
For any object of we have the following.
The functor is an additive right exact functor. Moreover, it commutes with small colimits.
The functors and are additive left exact functors. Moreover, commutes with limits and sends colimits to limits.
Proof.
Recall that direct sums are both products and coproducts. (Co)kernels are (co)equalisers. An additive functor is left (resp. right) exact if and only if it commutes with kernels (resp. cokernels). So we just need to prove that commutes with small colimits, commutes with small limits and sends colimits to limits. The fact that commutes with small colimits follows from the fact that it is left adjoint. Similarly we prove that commutes with limits. Let us prove that sends colimits to limits. Note that small colimits in are conical colimits in the enriched sense (2.29). Then the result follows from the formula (2.28). ∎
3.2. Category of modules
Let be a small -category. A -module is a -functor A morphism of modules is a -natural transformation. The category of -modules is denoted by
A morphism in this category is a -natural transformation (2.7). The set of morphisms will be denoted by
For an object of we consider the evaluation functor
Proposition 3.2.
The category of -modules is bicomplete. Moreover, for any object the evaluation functor preserves limits and colimits. In particular, for any functor from a small category there isomorphisms
which are natural in and and the following diagrams are commutative.
Proof.
Theorem 3.3.
For a small -category the category of -modules is an abelian category satisfying Grothendieck axioms and
Proof.
Let be morphisms of -modules. Consider a collection indexed by objects of Let us check that is a morphism of -modules using the definition (2.7). Since and are additive functors (Proposition 3.1), they commute with differences of morphisms. Combining this with the fact that composition in is bilinear, we obtain that is a morphism of -modules. Therefore is a subgroup. In particular, has a natural structure of a group. Using that the composition in is defined component-wise and the composition in is bilinear, we obtain that it is bilinear in Therefore, is a preadditive category.
Proposition 3.2 implies that has finite products and coproducts. Consider a zero -module defined by Using the definition of a morphism of -modules (2.7), it is easy to check that is a zero object in Since any preadditive category with finite products is additive [popescu1973abelian, §2.1], we obtain that is an additive category. Moreover, since is bicomplete, we obtain that it is also preabelian.
In order to prove that is abelian, we need to prove that for any morphism the morphism is an isomorphism. Proposition 3.2 imply that there are isomorphisms and which are natural in the sense that the diagram
is commutative. Using that the morphism is an isomorphism in we obtain that is an isomorphism. By Lemma 2.1 we obtain that the morphism is an isomorphism. Therefore is abelian.
Let us finally prove that satisfies the axiom and the axiom can be proved similarly. Consider a collection of short exact sequences We need to prove that the sequence is short exact. Since kernels, cokernels and images are defined component-wise in we know that is short exact for any and , and we need to prove that is short exact. By Proposition 3.2 we have natural isomorphisms and such that the diagram
is commutative in . Combining this with the fact that satisfies we obtain that the sequence is short exact. ∎
3.3. Induced modules
Let be a -functor between small categories. Then we consider functors
defined by the formulas and (see (2.34)). Then there are adjunctions
The -module will be called the restricted module of the -module is called the induced module of , and is called the coinduced module of By (2.41) we have the following explicit formulas
If we have two -functors and we have [kelly1982basic, Th. 4.47]
Projective (resp. injective) objects of the abelian category of -modules are called projective (resp. injective) -modules.
Proposition 3.4.
Let be a -functor between small -categories. Then the following holds.
is an exact functor, is right exact functor, is left exact functor.
sends projective -modules to projective -modules.
sends injective -modules to injective -modules.
Proof.
(1). It follows from (3.9) and the fact that left adjoint functor is left exact and right adjoint functor is left exact.
(2). Let be a projective -module. Then the functor is exact. Since is also exact, we obtain that is exact. Hence is projective.
(3). Similar. ∎
Proposition 3.5.
If is surjective on objects, then is a faithful functor. Moreover, in this case for any -module the counit of adjunction is an epimorphism, and the unit of adjunction is a monomorphism.
Proof.
For two morphisms of -modules the equation implies for any object of Since is surjective on objects the equation for each implies for each It follows that is faithfull.
Using that is left adjoint to we obtain that the composition
is identical, where is the unit of the adjunction, and is the counit of the adjunction. Therefore is an epimorphism. Since is faithfull, we obtain that is also epimorphism. Similarly we prove that is a monomorphism. ∎
3.4. Free modules
Any set can be treated as a discrete -category, whose objects are elements of and the hom-sets are defined so that and for Then is isomorphic to the category of families of objects from We will identify -modules with these families.
For any small -category we treat as a discrete -category. Then we have the canonical embedding of -categories So we can consider the functors
defined as So the forgetful functor sends a -module to the family and and are left and right adjoint functors to the forgetful functor
Equations (3.10) imply there are the following explicit formulas
A -module of the form is called a free module, and the -module of the form is called cofree -module.
Proposition 3.5 implies that for any -module we have an epimorphism from the free -module, and a monomorphism into a cofree -module
Proposition 3.6.
Assume that has enough of projectives. Then has enough of projectives. Moreover, a -module is projective if and only if it is a direct summand of a module of the form such that is a projective object of for any
Proof.
Consider a -module and apply the forgetful functor Then for each there exists an epimorphism from a projective module Therefore, we have an epimorphism in the category of -modules Since is right exact, the morphism is an epimorphism. Composing with the epimorphism (3.16), we obtain an epimorphism Proposition 3.4 implies that is projective. Then has enough of projectives. Note that, if is a projective -module, then the epimorphism splits, and is a direct summand of ∎
Similarly we can prove the following.
Proposition 3.7.
Assume that has enough of injectives. Then has enough of injectives. Moreover, a -module is injective if and only if it is a direct summand of a module such that is an injective object of for any
3.5. Flat and projective categories
We say that an object of is -flat (resp -projective), if (resp. ) is an exact functor. We say that a -category is -flat (resp. -projective), if is -flat (resp. -projective) for any objects
Lemma 3.8.
If is -flat (resp. -projective), then (resp. ) is an exact functor.
Proof.
It follows from the formulas (3.15) and axioms ∎
Lemma 3.9.
Let and be sets and be a map. If we treat and as discrete -categories, and as an -functor, then the functors can be described by the formulas
Proof.
It follows from (3.10). ∎
Proposition 3.10.
Let be a -functor between small -flat -categories. Assume that has enough of projectives. Then for any -module the -module is -acyclic.
Proof.
Denote by a projective resolution of in Using Lemma 3.8 and Proposition 3.6 we obtain that is a projective resolution of Then is equal to the homology of the chain complex The commutativity of the diagram
and equations (3.11) imply that Lemma 3.9 implies that the functor is exact. Lemma 3.8 implies that is exact. Therefore is also exact. Hence for ∎
Similarly we obtain the following.
Proposition 3.11.
Let be a -functor between small -projective -categories. Assume that has enough of injectives. Then for any -module the -module is -acyclic.
3.6. Bar resolution
For a small -category we consider the adjunction (3.14)
For a -module we have The unit of the adjunction is defined as the composition
The counit
has components, which are adjoint to
Consider the associated comonad on the category of -modules
The comultiplication is defined by the whiskering
We denote by the bar construction associated with this comonad [ivanov2021homology, §1], [appelgate1969homology]. So is a simplicial -module such that By induction we obtain
The face maps are defined by the formula and the degeneracy maps are defined by
Let us give a more detailed description of the face maps The face map is induced by the evaluation map For the face map is induced by the map
And similarly, the map is induced by the map
For any simplicial object in an abelian category, we denote by its non-normalised chain complex, and by its normalised chain complex. In particular, we consider the chain complexes and We will also use notation
Proposition 3.12.
Let be a -functor between small -flat -categories. Assume that has enough of projectives. Then the chain complex of the bar construction is an -acyclic resolution of In particular, the left derived functors of can be described as
Proof.
By [ivanov2021homology, Lemma 1.2] the augmented simplicial object has an extra degeneracy map. Therefore the chain complex is homotopy equivalent to Since is faithful and exact, and commutes with any additive functor, we obtain that for and The fact that is -acyclic follows from Proposition 3.10. ∎
Dually, we consider the adjunction
where The counit of the adjunction is defined as the composition
The unit of the adjunction
is induced by the morphisms adjoint to The corresponding monad is defined by the formula
Consider the corresponding cobar cosimplicial endofunctor on such that Using the isomorphism
by induction we obtain
The coface maps for are defined as follows. For they are dual to the morphisms induced by the composition (3.25), (3.24). The coface map is induced by the unit (3.29) The codegenerasy maps are induced by the counit (3.28).
Proposition 3.13.
Let be a -functor between small -projective -categories. Assume that has enough of injectives. Then the chain complex of the cobar construction is an -acyclic resolution of In particular, the right derived functors of can be described as
Proof.
Similar to Proposition 3.12. ∎
3.7. Homology of augmented categories
Assume that is an augmented -category with the augmentation (Subsection 2.8). Note that Therefore the augmentation induces functors of the trivial module, invariants and coinvariants
defined by the formulas
Then there are adjunctions
and Proposition 3.5 implies that is a faithfull functor.
If we expand the definition of the functor of trivial module, we obtain that for we have
and the morphism is defined as the composition
By (3.10) the functor of coinvariants can be described as the coequaliser
and the functor of invariants as the equaliser
The commutativity of the diagram
and Lemma 3.9 imply the isomorphisms
Example 3.14.
If is the category of abelian groups, is a group and is the group algebra, then -modules are -modules in the usual sense, and and the functor of invariants and coinvariants coincide with the usual functors
Assume that has enough of projectives. We define homology of an augmented -category as
Homology with the trivial coefficients are denoted by
Similarly, if has enough of injectives, we define cohomology as
Consider a simplicial object in whose components are
The face maps are defined so that for they are induced by compositions
The morphism is induced by the map adjoint to The morphism is induced by the augmentation The degeneracy maps are defined induced by the identity map
Theorem 3.15.
Let be an augmented -flat -category. Assume that has enough of projectives. Then for any -module there is an isomorphism
Proof.
Consider the nerve of (Subsection 2.8). Note that
Corollary 3.16.
Under the assumption of Theorem 3.15 there is an isomorphism
Consider a cosimplicial object in whose components are
The coface maps for are described as follows. For they are dual to the compositions (3.48). The morphism is dual to the map induced by the augmentation The coface map is induced by the unit (3.29) The codegenerasy maps are induced by the counit (3.28).
Theorem 3.17.
Let be an augmented -projective -category. Assume that is -projective and has enough of injectives. Then
Proof.
Similar to Theorem 3.15. ∎
Note that
Therefore, we obtain a corollary.
Corollary 3.18.
Under the assumption of Theorem 3.17 we have an isomorphism
4. Magnitude homology
4.1. Magnitude homology of an enriched category
Let us recall the definition of the magnitude homology from [leinster2021magnitude] (see also [hepworth2022magnitude]). We will give a slightly more general definition, which is more natural for our setting. We don’t assume that the base monoidal category is semicartesian. Instead of this we assume that the -enriched category is augmented (Subsection 2.8). This is a generalisation because any category enriched over a semicartesian category has a unique augmentation.
Let be a symmetric monoidal category, and be an abelian cocomplete closed symmetric monoidal category. Assume that is a strong monoidal functor. For an augmented -category we consider a simplicial object in whose components are
The face maps are defined for as follows. For they are induced by compositions
For they are induced by the augmentation
We denote by the normalised complex associated with the simplicial object Then the magnitude homology is the homology of this complex
If is a Bénabou cosmos, and is an abelian Bénabou cosmos, we can consider the -category defined in Subsection 2.7. Using that is strong monoidal, we obtain that the morphism is an isomorphism. Therefore and we obtain an augmentation
Theorem 4.1.
Let be a Bénabou cosmos, be an abelian Bénabou cosmos, be a strong monoidal functor and be an augmented -category. Assume that has enough of projectives and is -flat for any objects of . Then the category of -modules has enough of projectives and the magnitude homology is isomorphic to the homology of the augmented -category
Proof.
Let us recall the definition of the magnitude cohomology [hepworth2022magnitude]. It is defined in the same setting as homology. We consider a cosimplicial object of defined by the formula
and the associated normalised cochain complex The magnitude cohomology is defined as the cohomology of this cochain complex
Theorem 4.2.
Let be a Bénabou cosmos, be an abelian Bénabou cosmos, be a strong monoidal functor and be an augmented -category. Assume that has enough of injectives, is -projective for any objects of and is -projective. Then the category of -modules has enough of injectives and the magnitude cohomology is isomorphic to the cohomology of the augmented -category
Proof.
Remark 4.3 (Magnitude homology with coefficients).
4.2. Distance modules over a quasimetric space
A generalized metric space is a set equipped with a distance map such that
for any A quasimetric space is a generalized metric space such that
The following theory could be developed for any generalized metric space, but for the case of a quasimetric space there are several simplifications. For example, the definition of invariants and coinvariants of distance modules for the general case of generalized metric spaces is more complicated. So, for simplicity, further we will assume that is a quasimetric space. Elements of are called points. We say that a point lies between points and denoted by
if
In this section we use the abelian Bénabou cosmos of -graded modules over a commutative ring (for details see Subsection 7.1 in Appendix). A distance module over a quasimetric space consists of:
a family of -graded modules indexed by points of
morphisms of degree for each pair of points such that
satisfying conditions: and
An element will be called an element of over the point of degree . For such an element and any point such that we set
So any element of over defines an element over of degree
Then the equation (4.14) can be rewritten as
A morphism of distance modules is a family of morphisms of -graded modules of degree such that for any element of over a point and any point such that we have
The category of distance modules over is denoted by
For an -graded module we denote by the distance module such that
and for any This construction defines a functor
For a distance module we define the -graded module of invariants as follows
where
Dually we define coinvariants
where
These two constructions define functors
Further we assume that is a subgroup such that for any points (for example, if is a digraph, we can take ). Then we consider a full subcategory
that consists of distance modules such that for any Such distance modules will be called -graded. Then the functors of invariants and coinvariants of distance modules, and the functor of trivial distance module can be restricted to the functors
Denote by an augmented category enriched over whose objects are points of and hom-objects are defined as
If we denote by the generator of then and the composition is defined so that
The augmentation is defined by the formula
Proposition 4.4.
The category of -modules is isomorphic to the category of -graded distance modules
and the functors of trivial distance module, invariants and coinvariants for distance modules defined by (4.19), (4.21) (4.23) correspond to the functors trivial module, invariants and coinvariants of modules defined by (3.35). In particular has enough of projectives and injectives and there are adjunctions
Proof.
The isomorphism of categories is obvious. The fact that the functors of trivial modules coincide follows easily from the formula (3.37).
Let us prove that coinvariants in the sense of distance modules defined by the formula (4.23) coincide with coinvariants in the sense of modules over enriched categories defined by the formula (3.38). The formula (3.38) implies that that we need to consider the cokernel of the difference of two morphisms
defined by the formulas and where
Note that both and send the summand with the index to the summand with the index Therefore we can present them as and for morphisms
Denote by their restrictions to the corresponding summand. Then and if we have and Therefore
It follows that
Therefore
Hence these two definitions of coinvariants coincide.
The proof for the case of invariants is similar. ∎
Further we consider the trivial magnitude
Proposition 4.5.
For a distance module there is an isomorphism
that sends to
Proof.
Note that by the definition of the trivial distance module for any element of over a point and we have Therefore for any we have Therefore On the other hand, it is easy to see that for any we can define a morphism by the formula The assertion follows. ∎
Remark 4.6 (Right vs. left distance modules).
The distance modules that we defined may be called right distance modules. Similarly one can define left distance modules so that we have morphisms of degree and the category of left distance modules over is equivalent to the category of -modules. Note that where is a quasimetric space with the same set of points and a quasimetric defined by Therefore, the category of left distance modules over is equivalent to the category of right distance modules over
4.3. Magnitude homology of a quasimetric space
Let be a quasimetric space. For a tuple of points in we set
Let us recall the definition of the magnitude homology of a quasimetric space. For a real number we consider a simplicial -module which is freely generated by tuples of points such that
Tuples with are identified with zero in Then the -th face map is the map of deletion of the -th point, and the -th degeneracy map is the map of doubling of -th point. The assotiated normalized complex is denoted by
Then the magnitude homology of is defined as the homology of this chain complex
Now we give a definition of magnitude homology with coefficients in a distance module We consider a simplicial module whose components are defined by the formula
Here the summation is taken over all tuples of points of such that For an element we denote by
the corresponding element of from the summand indexed by The face maps for are defined as follows. For we have
and for we have
The degeneracy maps are defined by
The corresponding normalised complex will be denoted by
The magnitude homology of with coefficients in is defined by
It is easy to see that
where
Theorem 4.7.
Let be a subgroup such that for any points of Then the category of -graded distance modules has enough of projectives, the functor of coinvariants
is right exact, and the magnitude homology is isomorphic to its left derived functors
Proof.
In order to prove this theorem we will use Theorem 4.1. We consider the Bénabou cosmos of non-negative real numbers with infinity and abelian Bénabou cosmos of modules graded by real numbers. The strong monoidal functor is defined so that for and It is a strong monoidal functor with the obvious isomorphism Also note that (7.6) implies that the -graded modules are -flat and projective. So the functor sends objects of to -flat and -projective objects. The quasimetric space is treated as a -category in the usual way. The augmentation is uniquely defined. Then is the -category defined by the formulas (4.29), (4.30), (4.31). So the statement follows from Proposition 4.4 and Theorem 4.1. ∎
4.4. Magnitude cohomology of a quasimetric space
Let us recall the definition of the magnitude cohomology of from [hepworth2022magnitude]. For each real we consider the cosimplicial module
and the associated normalized cochain complex
whose cohomology is the magnitude cohomology of
Now we define the magnitude cohomology of with coefficients in a distance module . For a real we consider a cosimplicial module such that
For we denote by its projection in Then the face maps for are defined as follows. For we have
and for we have
and
The codegeneracy maps are defined by The associated normalized cochain complex is denoted by
The magnitude cohomology of with coefficients in is defined as cohomology of this complex
It is easy to see that
Theorem 4.8.
The category of -graded distance modules has enough of injectives, the functor of invariants
is left exact, and the magnitude cohomology is isomorphic to its right derived functors
Proof.
The proof is similar to the proof of Theorem 4.7. ∎
Remark 4.9.
Note that in the formula (4.67) we have in the subscript of the right-hand term. For this reason, we write in superscript in the notation for magnitude cohomology .
Corollary 4.10.
For any -graded distance module there magnitude cohomology with coefficients in can be presented as the Ext in the category of distance modules
5. Magnitude homology of a finite quasimetric space
5.1. Distance algebra of a finite quasimetric space
Let be a finite quasimetric space (§4.2) and be a subgroup of the additive group of real numbers such that for any points of Consider a -graded algebra which as a -module is freely generated by pairs of points such that and the product is defined by
The algebra will be referred to as the distance algebra. Note that the elements are orthogonal idempotents and is the unit of this algebra. So the set is a complete set of orthogonal idempotents. Note that in order to define the unit, we need to use that is finite. The -grading is defined by the formula In other words, is spanned by pairs such that In particular, is spanned by the idempotents
Proposition 5.1.
There is an equivalence of categories
that sends a -graded distance module to a -graded right -module
where the structure of -module is defined by if and by if
Proof.
It is easy to check that the construction from the statement defines a functor . Let us construct the functor in the opposite direction. If is a graded right -module, we construct a distance module so that and for The fact that is a complete set of orthogonal idempotents implies that Using this, the proof that is a functor opposite to the functor is straightforward. ∎
5.2. Magnitude homology of a finite quasimetric space
For a finite quasimetric space we denote by the homogeneous ideal of generated by all pairs of positive degree. It is easy to check that if is the cardinality of then
Therefore, is a nilpotent ideal. We set
As an algebra is isomorphic to the product where each factor is generated by the corresponding idempotent If is a principal ideal domain, then the Jacobson radical of is trivial. Using that is nilpotent, in this case we obtain that is the Jacobson radical of (that is why this notation was chosen).
Since is a quotient graded algebra of , it can be treated as a graded right -module and a graded left -module. This graded module is concentrated in degree zero and it is equal to the direct sum of graded submodules
where Using this formula we obtain that corresponds to the trivial distance module
Proposition 5.2.
If is a -graded distance module over and is the corresponding graded -module, then
Proof.
Since we have Since we obtain that is a direct sum of -modules
Comparing this with the formula (4.24), we obtain the isomorphism The second isomorphism can be proved similarly. ∎
In the statement of the next theorem we use bigraded Tor and Ext functors. For the definition see Subsection 7.2 of the Appendix.
Theorem 5.3.
Let be a commutative ring, be a subgroup, be a finite quasimetric space such that and be a -graded distance module over . Then for any there are isomorphisms
In particular, we obtain
Remark 5.4 (Bar resolution of ).
The statement of Theorem 5.3 can be extended by constructing an explicit projective resolution of the graded right -module . The resolution from Proposition 3.12 together with the equivalence from Proposition 5.1 defines a graded projective resolution of the graded -module such that
and the differential is defined by where is defined by
The structure of -module on is defined by if and and by otherwise. The module is projective, because there is an isomorphism of -modules
where is the projective summand of corresponding to the idempotent
Remark 5.5.
Apparently, the isomorphism could be generalized to the case of any quasimetric space. In order to do this, for a right distance module over and a left distance module over we need to introduce the tensor product which is an -graded module. Then we need to define the Tor functor of distance modules as the derived functor and show that
5.3. Magnitude cohomology as a Yoneda algebra
Usually the Yoneda product is defined in terms of -fold extensions [maclane2012homology, Ch.III,§5] [buchsbaum1959note], [yoneda1958note], [yoneda1954homology] but we use the description in terms of projective resolutions that can be found in [maclane2012homology, §6 Exercise 2].
Let be an abelian category with enough of projectives and are its objects. Then the Yoneda product
is defined as follows. Let be a projective resolution of and be a projective resolution of Assume that and are two cycles representing elements from and The fact that is a cocycle means that Therefore, induces a morphism This morphism can be lifted to the morphism of resolutions uniquely up to homotopy.
Then the Yoneda product is defined as
The Yoneda product is bilinear. This allows us to define a structure of a graded algebra on which is called Yoneda algebra of
If is an abelian group, is a -graded algebra and is a -graded -module, then the -graded Yoneda algebra of is defined as
where the product is also defined by the Yoneda product (see (7.19)). If is a projective resolution of then is a projective resolution of An element of can be presented by a homomorphism such that Then for another such a homomorphism we have
All these definitions work similarly for left and right modules.
Now let us recall the definition of the product for the magnitude cohomology of a quasimetric space from [hepworth2022magnitude]. First we define the structure of a dg-algebra on
If and then is defined by
if and and by
otherwise. This product induces a product on the magnitude cohomology that defines a structure of a -graded algebra on
Theorem 5.6.
Let be a commutative ring, be a subgroup and be a finite quasimetric space such that Then the -graded magnitude cohomology algebra of is isomorphic to the -graded Yoneda algebra of the -graded left -module
Remark 5.7.
In Theorem 5.6 we treat as a left module. If we treated as a right module, a similar proof would give an isomorphism where “op” means that we consider the opposite product defined by for and In order to avoid this construction of opposite algebra in the statement of Theorem 5.6 and to simplify the proof, we decided to formulate it in terms of left modules.
Proof of Theorem 5.6.
Set In this proof we deal with left -graded -modules. We need a left-module version of the projective resolution of from Remark 5.4. Its components are defined by
and the differential is defined by where is given by
The structure of left -module on is defined by if and and by otherwise. There is an isomorphism of -modules
Note that
Here is generated by the projection It follows that for any Using the isomorphism (5.27) we obtain
where a map corresponds to a map defined by
Here we assume that
It is easy to check that this isomorphism is compatible with the differentials, and we have an isomorphism of complexes
which induces an isomorphism for cohomology
Let us prove that this isomorphism is an isomorphism of bigraded algebras.
Assume that we have an element in the magnitude cohomology Take a representing cocycle Then is also a cocycle and it induces a homomorphism Consider a homomorphism defined by the formula
It is easy to check that is a homomorphism of right -modules such that the diagram
is commutative. Further for any we define by the formula
We claim that defines a morphism of chain complexes:
If we prove this, the statement of the theorem follows easily, because if is another cocycle, then it is easy to check that
So the rest of the proof is devoted to the proof of the equation (5.37).
If then the statement is obvious, so we will assume The differentials are defined by the formulas
First we note that
Indeed, both maps applied to give if and otherwise (here we assume ). On the other hand, we have
Then, using the fact that is a cycle, we obtain
Finally we note that
Indeed, if then both parts of the equation are zero, because and and if then both parts of the equation are equal to The equation (5.37) follows from the equations (5.40),(5.42),(5.43). ∎
6. Magnitude homology of a digraph
6.1. Background on representations of quivers
A quiver is a quadruple such that are sets and are maps. Elements of are called vertices, elements of are arrows, for an arrow we say that is the start of and is the target of A quiver is called finite, if are finite.
Let be a commutative ring. A representation of a quiver is a collection of -modules indexed by vertices and a collection of homomorphisms for each arrow where and A morphism of representations is a collection of homomorphisms indexed by vertices such that for any arrow from to The category of representations will be denoted by
A path of length in a quiver is a sequence of arrows such that With each vertex we also associate a trivial path of length zero. For a finite quiver we define the path algebra as a -algebra, whose -basis consists of all paths and trivial paths. The product is defined by concatenation of paths, and it vanishes if the concatenation does not exist. The algebra has a unit defined as the sum of trivial paths Note that in order to define the unit we use that is finite. It is well-known that the category of representations of a finite quiver is equivalent to the category of right -modules
(see [auslender, Ch.III, Th.1.5], [crawley1992lectures, §1, Lemma], [hazewinkel2006algebras2, Theorem 2.4.1]). This equivalence sends a representation to a -module such that for any and any arrow such that and if The functor in the opposite direction sends a module to a representation defined by and
A quiver with relations is a pair where is a finite quiver and is an ideal of . We say that a representation of a quiver satisfies relations from if the annihilator of the corresponding -module contains The full subcategory of representations satisfying relations from is denoted by
The equivalence (6.1) can be restricted to an equivalence
(see [auslender, Ch.III, Prop.1.7], [assem2006elements, Ch.III, Th.1.6]).
All this theory can be developed in the graded setting. In this section we assume that and “graded” means “-graded”. A graded representation of a quiver is a collection of graded -modules indexed by vertices and a collection of homomorphisms of degree for each arrow where and A morphism of graded representations is a collection of homomorphisms of degree zero indexed by vertices such that for any arrow from to The category of graded representations will be denoted by The path algebra has a natural grading, where the -th homogeneous component consists of linear combinations of paths of length Similarly to the non-graded case we obtain an equivalence
If we assume that is a homogeneous ideal of , we can also similarly define a subcategory
such that there is an equivalence
Set Denote by the ideal of generated by all non-trivial paths. The ideal is called admissible if
for some (see [assem2006elements, Ch. II, Def. 2.1]). In this case is called bound quiver and is the bound quiver algebra (usually this terminology is used assuming is a field). In this case is finitely generated as a -module. The quotient
has a natural structure of a (right and left) -module. Note that is the direct sum of submodules
If is a field and is admissible, the bound quiver algebra is finite dimensional and is the Jacobson radical of (see [auslender, §III.1, Prop. 1.6], [assem2006elements, Ch.II, Lemma 2.1]). Moreover, in this case is a complete set of representatives of the isomorphism classes of the simple -modules and is the direct sum this set [assem2006elements, Ch. III, Lemma 2.1].
6.2. Distance modules as representations of quivers
We define a digraph as a quiver without loops and multiple arrows. In this case arrows can be identified with pairs of vertices So we can assume that and are projections. A finite digraph is a digraph such that is finite. Further in this section we assume that is a fixed finite digraph.
A path in the path algebra can be identified with a sequence of the vertices The trivial path is identified with A path is called shortest, if it is a shortest path from to A path which is not shortest will be called non-shortest. Denote by a homogeneous ideal generated by two types of relations
if and are two different shortest paths from to
if is a non-shortest path.
We say that a non-shortest path is a minimal non-shortest path, if and are shortest paths. It is easy to see that (R2) can be replaced by
(R2’)
if is a minimal non-shortest path.
The distance between vertices of as the infimum of lengths of paths from to So defines a finite quasimetric space with integral distances that we denote by the same letter . We will consider the category of -graded distance modules In this section all distance modules are -graded.
Proposition 6.1.
The ideal is admissible and there is an isomorphism of graded algebras
Moreover, there is an equivalence of categories
that sends a -graded distance module to a representation such that and for each
Proof.
If there are a least two shortest paths from one vertex to another, then these paths have lengths Non-shortest paths also have length Therefore There is a finite number of shortest paths in If we take which is larger than the length of any shortest path, then Hence is admissible.
As a corollary we obtain the following theorem.
Theorem 6.2.
Let be a commutative ring and be a finite digraph. Then there are isomorphisms
Moreover, the magnitude cohomology algebra is isomorphic to the Yoneda algebra of the left -module
7. Appendix. Graded modules and algebras
7.1. Abelian cosmos of graded modules
Let be a commutative ring and be an abelian group. In this subsection we will introduce the Bénabou cosmos of -graded -modules In this subsection the word “module” means “-module”.
An -graded module is a family of modules An element of will be called a homogeneous element of of degree The degree of a homogeneous element will be denoted by A morphism of -graded modules of degree is defined as a family of homomorphisms The degree of a morphism will be denoted by Therefore for a homogeneous element of we have The composition of such morphisms is defined in the obvious way so that We denote by the abelian group of homomorphisms of degree
and denote by the -graded module, whose components are
For two -graded modules and we denote by an -graded module such that
Then it is easy to check that for any -graded modules there is an isomorphism of -graded modules
For an -graded module and a we denote by the shifted -graded module
For any (non-graded) module we also denote by the -graded module concentrated in degree
It is easy to see that for any -graded module there are isomorphisms
The category of -graded modules is a category whose objects are -graded modules and morphisms are morphisms of degree
It is easy to see that the tensor product the inner hom and the unit make it an abelian Bénabou cosmos. One can say that it as a category it is a product of the categories of -modules Therefore, this category has enough of projectives and injectives.
7.2. Graded modules over graded algebras
By a -graded -algebra we mean a collection of -modules equipped with bilinear homomorphisms and an element satisfying the associativity and unit axioms (one can say that a graded algebra is a monoid in the cosmos of -graded modules). Traditionally people consider an ordinary algebra associated with the graded algebra, and call it graded algebra. We will also allow such freedom of speech, replacing by if it does not lead to confusion. However, the situation with tensor product and hom of -graded -modules is more complex, and there such freedom of speech can lead to confusion. So let us describe in details relations between hom-sets and tensor products of -graded -modules and -modules. Further we assume that the word “graded” means “-graded” for the fixed .
If is a graded algebra, a right -graded -module is a collection of -modules equipped with bilinear maps satisfying the exterior associativity and unit axioms. A left graded -module is defined similarly. If and are right graded -modules, then
is a graded -module whose -th component consists of collections of -homomorphisms such that for any The hom-set in the category of graded -modules is the -component of this graded module
So we can say that
Now assume that is a right graded -module and is a left graded -module. Then the tensor product
is defined as the graded -module, whose -th component is defined as the quotient of the -module by the submodule generated by elements of the form It is easy to see that
For a graded -module we denote by the corresponding -module. We will also consider a bigger -module
The elements of are families where Then for the product is defined by The same notations are used for graded -modules. Using that the tensor product over commutes with direct sums, it is easy to check that there is an isomorphism
However, for hom-sets the formula is more complicated
It follows from the isomorphism If both graded modules are non-trivial only in finitely many degrees, then products in the last formula can be replaced by direct sums.
The derived functors of the functors
are denoted by
where The formula (7.10) implies that the bigraded Ext can be presented as the ordinary Ext in the category of graded modules
Note that this implies the symmetric description of the bigraded ext
Using this, formulas (7.14), (7.15) and the fact that for a free graded resolution of the complex is a free graded resolution of , we obtain